Fill in the blanks with positive integers:

(3 + sqrt(6))*3 = ___ + ___ * sqrt(6)

Answer by ikleyn(52781)

.
Fill in the blanks with positive integers:

(3 + sqrt(6))*3 = ___ + ___ * sqrt(6)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


(3 + sqrt(6))*3 = 9 + 3*sqrt(6).         ANSWER